# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

### v3

@@ Abstract.
[…]
@@ 1. Introduction
Recent advancements in AI (Goodfellow et al., 2014; Li et al., 2022; Touvron et al., 2023; Alayrac et al., 2022) are propelled by a class of models called foundational models (FMs), including large language models (LLMs) and vision-language models (VLMs). FMs leverage billion-scale parameters for improved learning (Yang et al., 2024; Zhang et al., 2024) but impose substantial demands on memory, energy, and compute resources. Recent research has focused on various model compression techniques such as pruning (Frantar and Alistarh, 2023; Ashkboos et al., 2024a) and quantization (Lin et al., 2024a; Shao et al., 2023; Guo et al., 2023; Kang et al., 2024) to reduce memory footprint and computational costs, and make FM inference service more accessible to resource-constrained devices.
Model pruning (Frantar and Alistarh, 2023) reduces memory footprint by removing ineffectual model parameters, such as individual weights (unstructured) or blocks of weights (structured), and storing sparse tensors in a compressed format (Jeong et al., 2023). However, pruning of FMs may be infeasible due to, significant accuracy drops even at low pruning ratios (Kuzmin et al., 2024; Yin et al., [n. d.]) and potential demand for compute and memory-intensive fine-tuning to regain accuracy. Model quantization, on the other hand, reduces the size of a target model by representing weights and/or activations at low precision (Ramachandran et al., 2024b; Kang et al., 2024; Ramachandran et al., 2024a). Recent works on quantization (Dettmers et al., 2022; Guo et al., 2023) have identified that quantizing LLMs is considerably more challenging than quantizing traditional DNNs (Li et al., 2021; Tambe et al., 2020) due to the emergence of large magnitude features known as outliers (Zadeh et al., 2020). These outliers significantly impact model accuracy.
Table 1. Comparison of MicroScopiQ with prior outlier-aware co-design techniques in group A (Zadeh et al., 2020; Park et al., 2018) and B (Guo et al., 2023).
@@ Methods
Techniques in group A, such as OWQ (Lee et al., 2024), SpQR (Dettmers et al., 2023) (algorithm) and GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018), SDQ (Jeong et al., 2024) (architecture co-design) exhibit low accuracy degradation. This is because they typically store outliers at high precision separated from lower precision inliers. However, these techniques result in, (a) low compression factor with high effective bit-width (EBW [^1] The average number of bits used to represent each quantized parameter of a model.) and (b) inefficient hardware and unaligned memory access.
@@
On the other hand, techniques in group B, such as AWQ (Lin et al., 2024a) (algorithm) and OliVe (Guo et al., 2023) (architecture co-design) quantize outliers at the same precision as inliers following different strategies. AWQ tries to identify a separate outlier-specific scale factor via channel-wise scaling. OliVe uses the ”flint” data format (Guo et al., 2022) for inliers and ”abfloat” (Guo et al., 2023) for outliers, both at 4-bit precision. These techniques mitigate the unaligned memory access while providing high compression. However, they suffer from significant accuracy degradation, particularly at ultra-low bit widths. This may be attributed to the reduced representational range available to outliers at ultra low-precision. Additionally, these methods (Guo et al., 2023) rely on a specific kind of locality of presence for the outliers, that might not be true for all the FMs, as we shall demonstrate in this work (section 3.1).
@@
Based on the limitations of existing solutions discussed above, we identify that assignment of higher bit-width for outliers is essential for good accuracy while for aligned memory and hardware efficiency a consistent bit-budget 222number of bits per weight/activation element and data type per tensor element is desired. Here, by consistency we mean that on average each scalar should be represented by a fixed bit-width of a particular data-type. However, these demands are generally conflicting.
Contributions. To provide a unified solution, we investigate on a fundamental question: Can pruning be effectively leveraged to complement outlier-aware quantization in achieving high accuracy while maintaining hardware efficiency?
@@
Towards achieving this, we present a novel co-design technique for the post-training quantization (PTQ) of FMs, namely MicroScopiQ. Our approach effectively leverages pruning with outlier aware quantization to achieve both memory alignment and improved accuracy. To effectively perform this, for a layer we prune the least important weights based on the Hessian information, and redistribute the additional bits of the outlier weights in the pruned locations. This ensures memory alignment while allowing outlier weights to have higher precision. Additionally, to maintain low error we use the recently proposed MicroScaling (MX) FP data format (Rouhani et al., 2023; Darvish Rouhani et al., 2023) to quantize outlier weights compared to MX-INT inlier quantization. To efficiently support outliers in a different format with different bits at different locations, we present an intelligent NoC architecture called, Redistribution and Coordination NoC (ReCoN). It offers minimal overhead and high throughput outlier processing and reorganization. We then present an accelerator that leverages ReCoN with a simple, homogeneous INT-PE array. Further, we extend the accelerator to be generic enough to support multiple bit-precision (2/4-bit operations). As summarized in Table 1, MicroScopiQ blends the advantages of group A and B techniques while mitigating their specific drawbacks.
@@ 2.1. Model Quantization
A typical quantization (Ramachandran et al., 2024a; Li et al., 2021) process involves two steps: establishing the quantization parameters, given the quantization data format $(\tau)$ and bit-width $(b)$, and mapping the high-precision tensor to the quantized representation. For a typical symmetric quantization (Zhao et al., 2024) (zero-point is 0) of a tensor $X$, the scale factor ($s$) is given by,
@@ 3.1. Limitations of existing techniques
OliVe (Guo et al., 2023) proposes a scheme to ensure aligned memory access by quantizing inliers and outliers at the same precision (low EBW), but using different data formats. To enable differentiation between the inlier and outlier formats, it prunes the value adjacent to the outlier for use as an identifier (see Figure 3(c)). However, OliVe results in significant accuracy degradation, especially at low precision (see Figure 2(b)), due to: 1) sacrificing a number encoding from inliers for exclusive use as an identifier, reducing the number of representable values in the quantized range, and 2) the rigid assumption of outlier locality–that outliers are never adjacent to each other and only inliers are almost always adjacent to outliers (see section 3.2), leading to unintended outlier pruning. Furthermore, OliVe requires a fairly complex PE design incurring significant encoding/decoding overheads to convert the different formats into a unified processing format (exponent-integer pair). We show that despite quantizing outliers at higher-precision and in a different format, we ensure aligned memory access, simple PE design and minimal hardware overhead.
Figure 2. ((a) Layer-wise distribution of outliers and adjacent outliers as a percentage of total number of weights, (b) Quantization accuracy comparison between OliVe-W4A16 and MicroScopiQ-W2A16 on various benchmarks.
@@ 3.2. Adjacent Outliers Matter
Similar to prior works (Guo et al., 2023; Park et al., 2018), we leverage the $3\sigma$ rule (Pukelsheim, 1994) to categorize weights as outliers. We visually demonstrate the distribution of outliers and adjacent outliers  [^3] We define adjacent outliers as two contiguous outliers along the dot-product dimension (see row 2 of the LLM weight matrix in Figure 3). as a percentage of the total number of weights in a layer across different FMs in Figure 2(a). As the orange box-plot shows, outliers depict a maximum percentage of $\sim 5.1\%$. Outliers are prevalent in FMs, and preserving their values is crucial for maintaining quantized model accuracy. Importantly, from the green box-plots, we observe that modern FMs on average possess $>0.5\%$ adjacent outliers per layer, with some FM layers showing peaks of $>2\%$. This is in stark contrast to the models evaluated by OliVe, such as BERT (Devlin et al., 2018) and OPT (Zhang et al., 2022) which have $<0.04\%$ adjacent outliers (two orders of magnitude lower than FMs like LLaMA3 and LlaVa). This indicates that while pruning values adjacent to outliers could have been ideal for models like BERT (Devlin et al., 2018), it is sub-optimal for most modern FMs as it removes crucial outlier values, leading to higher accuracy degradation. This is evident from Figure 2(b) where OliVe has significant accuracy degradation at 4-bit quantization due to its assumption on outlier locality. Unlike OliVe, MicroScopiQ does not naively prune adjacent values; instead it leverages the Hessian information (Frantar et al., 2022) to identify the least important values to prune, ensuring outlier preservation. This directly translates to high quantized model accuracy and MicroScopiQ at 2-bit consistently outperforms OliVe across different FMs.
@@ 4.1. Preliminaries
We evaluate the second order derivative of Equation 3, namely the Hessian (Hassibi et al., 1993; LeCun et al., 1989) through Taylor series expansion (Hassibi et al., 1993). Note, the Hessian is the same for all rows depending only on input data and is given as, $\mathbb{H}=2\mathbb{X}\mathbb{X}^{T}$. Its inverse is, $\mathbb{H}^{-1}=(2\mathbb{X}\mathbb{X}^{T}+\lambda\mathbb{I})^{-1}$. We leverage the Hessian information to identify weights with small “saliencies” for pruning (L17 in Algo. 1). The number of inliers to prune is determined by the number of outliers, as detailed later.
@@
4.2. Inlier and Outlier Weight Quantization
In step 1, each row to be quantized is first divided into multiple non overlapping contiguous macro-blocks (MaBs) of size $B_{M}=128$. All inliers are quantized within a MaB and share the scale factor. Each MaB is then subdivided into multiple non-overlapping contiguous micro-blocks ($\mu$Bs) of size $B_{\mu}=8$ with sixteen $\mu$Bs forming a MaB. The outliers present in each $\mu$B shares same scale(s). As depicted in Figure 3(a), step 1, the quantization process begins by first identifying inliers and outliers in each MaB by using the $3\sigma$ rule. A shared 8-bit power-of-two scale factor ($2^{I_{sf}}$), following Equation 1 is calculated for all inliers in a MaB and the inliers are quantized to 2-bit or 4-bit, resulting in MX-INT-(2/4)128 quantization. Interestingly, we observe that the inlier scale factor in each MaB is always a negative power of two for all FMs under consideration. We leverage this observation to reduce outlier magnitude, by multiplying all outlier values in a MaB with the inlier scale factor ($2^{I_{sf}}$) (can also be perceived as division by $2^{-I_{sf}}$, for conformity with Equation 2). This pre-processing helps make outlier quantization easier, by pre-reducing its dynamic range.
Unlike inliers, outliers are quantized per $\mu$B, to reduce quantization error due to shared scaling over a larger group size (see section 6). After identifying outliers present in a $\mu$B, we compute a shared 8-bit MXScale that is calculated by concatenating the level-1 power-of-two scale factor ($2^{O^{l_{1}}_{sf}}$) and level-2 microExponent ($\mu X$). The level-1 scale factor is calculated by following Equation 1 to obtain 7 or 5-bit MSBs of MXScale depending on size of $\mu X$ being 1 or 3-bit of the LSBs–corresponds to exponent size of the FP format (depicted in step 2 in Figure 3(a)). The outliers in a $\mu$B are scaled by ($2^{O^{l_{1}}_{sf}}$), following Equation 2 and then quantized to either e1m2/e3m4 FP-format for $b$ = 4 or 8-bit, respectively. Post quantization of outliers, the level-two scale factor or the $\mu X$ is obtained by extracting the common exponent among all outliers in a $\mu$B. This process results in a MX-FP-$b$8,8 outlier quantization. The final outlier scale factor is $2^{O_{sf}}$ where, $O_{sf}$ is expressed as ${O_{sf}}={O^{l_{1}}_{sf}+\mu X-I_{sf}}$. The term $I_{sf}$ in the final outlier scale factor accounts for multiplication by $2^{I_{sf}}$ (or division by the inverse) during outlier pre-processing.

### v4

@@ Abstract.
Quantization of foundational models (FMs) is significantly more challenging than traditional DNNs due to the emergence of large magnitude values called outliers. Existing outlier-aware algorithm-architecture co-design techniques either use mixed-precision, retaining outliers at high precision but compromise hardware efficiency, or quantize inliers and outliers at the same precision, improving hardware efficiency at the cost of accuracy. To address this mutual exclusivity, we propose MicroScopiQ, a novel co-design technique that leverages pruning to complement outlier-aware quantization. MicroScopiQ retains outliers at higher precision while pruning a certain fraction of least important weights to distribute the additional outlier bits; ensuring high accuracy, aligned memory and hardware efficiency. We design a high-throughput, low overhead accelerator architecture composed of multi-precision INT processing elements and a network-on-chip called ReCoN that efficiently abstracts the complexity of supporting high-precision outliers. Additionally, unlike prior techniques, MicroScopiQ does not assume any locality of outlier weights, enabling applicability to a broad range of FMs. Extensive experiments across diverse quantization settings demonstrate that MicroScopiQ achieves state-of-the-art quantization accuracy, while delivering up to $\mathbf{3}\times$ faster inference and $\mathbf{2}\times$ lower energy consumption compared to existing alternatives. Code is available at: MicroScopiQ-LLM-Quantization.git
Foundational Models, Quantization, Pruning, Hardware Accelerator
journalyear: 2025copyright: ccconference: Proceedings of the 52nd Annual International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japanbooktitle: Proceedings of the 52nd Annual International Symposium on Computer Architecture (ISCA ’25), June 21–25, 2025, Tokyo, Japandoi: 10.1145/3695053.3730989isbn: 979-8-4007-1261-6/2025/06ccs: Computer systems organization Systolic arraysccs: Computer systems organization Neural networksccs: Computer systems organization Data flow architecturesccs: Networks NoC
@@ 1. Introduction
Recent advancements in AI (Goodfellow et al., 2014; Li et al., 2022; Touvron et al., 2023; Chen et al., 2024) have been propelled by a class of models called foundational models (FMs), which encompass large language models (LLMs) and vision-language models (VLMs). FMs leverage billion-scale parameters for improved learning (Yang et al., 2024; Zhang et al., 2024) but impose substantial demands on memory, energy, and compute resources. Recent research has focused on various model compression techniques such as pruning (Frantar and Alistarh, 2023; Ashkboos et al., 2024a; Yin et al., 2023) and quantization (Lin et al., 2024a; Shao et al., 2023; Ramachandran et al., 2024b) to reduce memory and computational overhead, enabling efficient FM inference on resource-constrained devices.
Model pruning (Frantar and Alistarh, 2023) reduces memory footprint by removing ineffectual model parameters, such as individual weights (unstructured) or blocks of weights (structured), and storing sparse tensors in a compressed format (Jeong et al., 2023). However, pruning of FMs may be infeasible due to, significant accuracy drops even at low pruning ratios (Kuzmin et al., 2024; Yin et al., 2024) and potential demand for compute and memory-intensive fine-tuning to regain accuracy. Model quantization, on the other hand, reduces the size of a target model by representing weights and/or activations at low precision (Ramachandran et al., 2024b; Kang et al., 2024; Ramachandran et al., 2024a; Kundu et al., 2025). Recent works on quantization (Dettmers et al., 2022; Guo et al., 2023; Ramachandran et al., 2025b) have identified that quantizing LLMs is considerably more challenging than quantizing traditional DNNs (Li et al., 2021; Tambe et al., 2020) due to the emergence of large magnitude features known as outliers (Zadeh et al., 2020). These outliers significantly impact model accuracy and require specialized handling (Guo et al., 2023) compared to inliers.
Table 1. MicroScopiQ vs. prior outlier-aware quantization techniques, categorized into two groups, A (Zadeh et al., 2020; Park et al., 2018), B (Guo et al., 2023).
@@ Methods
Techniques in group A, such as OWQ (Lee et al., 2024), SpQR (Dettmers et al., 2023), SDQ (Jeong et al., 2024) (algorithm) and GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018)(architecture co-design) exhibit low accuracy degradation. This is because, they typically store outliers at high precision separated from lower precision inliers. However, these techniques result in, (a) low compression factor with high effective bit-width (EBW [^1] The average number of bits used to represent each quantized parameter of a model.) and (b) inefficient hardware and unaligned memory access.
@@
On the other hand, techniques in group B, such as AWQ (Lin et al., 2024a) (algorithm) and OliVe (Guo et al., 2023) (architecture co-design) quantize outliers at the same precision as inliers following different strategies. AWQ tries to identify a separate outlier-specific scale factor via channel-wise scaling. OliVe uses the “flint” data format (Guo et al., 2022) for inliers and “abfloat” (Guo et al., 2023) for outliers, both at 4-bit precision. These techniques mitigate the unaligned memory access while providing high compression. However, they suffer from significant accuracy degradation, particularly at ultra-low bit widths. This may be attributed to the reduced representational range available to outliers at ultra low-precision. Additionally, these methods (Guo et al., 2023) rely on a specific kind of locality of presence for outliers, that might not be true for all FMs, as we shall demonstrate in this work (section 3.1).
@@
Based on the shortcomings of existing solutions discussed above, we identify that assignment of higher bit-width for outliers is essential for good accuracy while for aligned memory and hardware efficiency a consistent bit-budget and data type per tensor element is desired. Here, by consistency we mean that on average each scalar within a tensor should be represented by a fixed bit-width of a particular data-type. However, these demands are conflicting.
Contributions. To provide a unified solution, we investigate on a fundamental question:
[…]
@@
Towards achieving this, we present a novel co-design technique for the post-training quantization (PTQ) of FMs, namely MicroScopiQ. Our approach effectively leverages pruning with outlier aware quantization to achieve both memory alignment and improved accuracy. To effectively perform this for a layer, we quantize outliers at twice the precision of inliers and prune the least important weights based on the Hessian information. We then redistribute the additional bits of the outlier weights in these pruned locations. This ensures memory alignment while allowing outlier weights to have higher precision. Additionally, to reduce error we use the recently proposed MicroScaling (MX) FP data format (Rouhani et al., 2023; Darvish Rouhani et al., 2023) to quantize outlier weights as opposed to MX-INT inlier quantization. While prior work such as SDQ (Jeong et al., 2024) also combined pruning and quantization, it contrasts with our approach in its limited outlier flexibility, lower compression factor, and unaligned memory access.
To efficiently support outliers in a different format with different bits at different locations in hardware, we present an intelligent NoC architecture called, Redistribution and Coordination NoC (ReCoN). It offers minimal overhead and high throughput outlier processing and reorganization. We then present an accelerator that leverages ReCoN with a simple, homogeneous INT-PE array. Additionally, we extend the accelerator to be generic enough to support multiple bit-precision (2/4-bit operations). As summarized in Table 1, MicroScopiQ blends the advantages of group A and B techniques while mitigating their specific drawbacks.
@@ 2.1. Model Quantization
A typical quantization (Ramachandran et al., 2024a; Li et al., 2021; Ramachandran et al., 2025b) process involves two steps: establishing quantization parameters given the quantization data format $(\tau)$ and bit-width $(b)$, and mapping the high-precision tensor to the quantized representation. For a typical symmetric quantization (Zhao et al., 2024) (zero-point is 0) of a tensor $X$, the scale factor ($s$) is given by,
@@ 2.2. Microscaling Data Format
Figure 2. (a) Layer-wise distribution of outliers and adjacent outliers as a percentage of total number of weights, (b) Quantization accuracy comparison between OliVe-W4A16 and MicroScopiQ-W2A16 on various benchmarks.
@@ 3.1. Limitations of existing techniques
OliVe (Guo et al., 2023) proposes a scheme to ensure aligned memory access by quantizing inliers and outliers at the same precision (low EBW), but using different data formats. To enable differentiation between the inlier and outlier formats, it prunes the value adjacent to the outlier for use as an identifier (see Figure 3(c)). However, OliVe results in significant accuracy degradation, especially at low precision (see Figure 2(b)), due to: 1) sacrificing a number encoding from inliers for exclusive use as an identifier, reducing the number of representable values in the quantized range, and 2) the rigid assumption of outlier locality–that outliers are never adjacent to each other and only inliers are almost always adjacent to outliers (see section 3.2), leading to unintended outlier pruning. Furthermore, OliVe requires a fairly complex PE design incurring significant encoding/decoding overheads to convert the different formats into a unified processing format (exponent-integer pair). In this paper, we show that despite quantizing outliers at higher-precision and in a different format, we ensure aligned memory access, simple PE design and minimal hardware overhead.
@@ 3.2. Adjacent Outliers Matter
Similar to prior works (Guo et al., 2023; Park et al., 2018), we leverage the $3\sigma$ rule (Pukelsheim, 1994) to categorize weights as outliers. We visually demonstrate the distribution of outliers and adjacent outliers  [^2] We define adjacent outliers as two contiguous outliers along the dot-product dimension (see row 2 of the LLM weight matrix in the center of Figure 3). as a percentage of the total number of weights in a layer across different FMs in Figure 2(a). As the orange box-plot shows, outliers depict a maximum percentage of $\sim 5.1\%$. Outliers are prevalent in FMs, and preserving their values is crucial for maintaining quantized model accuracy. Importantly, from the green box-plots, we observe that modern FMs on average possess $>0.5\%$ adjacent outliers per layer, with some FM layers showing peaks of $>2\%$. This is in stark contrast to the models evaluated by OliVe, such as BERT (Devlin et al., 2018) and OPT (Zhang et al., 2022) which have $<0.04\%$ adjacent outliers (two orders of magnitude lower than FMs like LLaMA3 and LlaVa). This indicates that while pruning values adjacent to outliers could have been ideal for models like BERT (Devlin et al., 2018), it is sub-optimal for most modern FMs as it removes crucial outlier values, leading to higher accuracy degradation. This is evident from Figure 2(b) where OliVe has significant accuracy degradation at 4-bit quantization due to its assumption on outlier locality. Unlike OliVe, MicroScopiQ does not naively prune adjacent values; instead it leverages the Hessian information (Frantar et al., 2022) to identify the least important values to prune, ensuring outlier preservation. This directly translates to high quantized model accuracy and MicroScopiQ at 2-bit consistently outperforms OliVe across different FMs.
@@ 4.1. Preliminaries
We evaluate the second order derivative of Equation 3, namely the Hessian (Hassibi et al., 1993; LeCun et al., 1989) through Taylor series expansion (Hassibi et al., 1993). Note, the Hessian is the same for all rows due to its dependence only on input data and is given as, $\mathbb{H}=2\mathbb{X}\mathbb{X}^{T}$. Its inverse is, $\mathbb{H}^{-1}=(2\mathbb{X}\mathbb{X}^{T}+\lambda\mathbb{I})^{-1}$. We leverage the Hessian information to identify weights with small “saliencies” for pruning (L17 in Algo. 1). The number of inliers to prune is determined by the number of outliers, as detailed later.
@@ 41      $\mathbb{W}_{(i+rB):,:}=\mathbb{W}_{(i+rB):,:}-\mathbb{E}\cdot\mathbb{H}^{-1}_{(i+rB):,i:(i+rB)}$
4.2. Inlier and Outlier Weight Quantization
In Step 1 (Figure 3(a), Algorithm 1), each row to be quantized is first divided into multiple non overlapping contiguous macro-blocks (MaBs) of size $B_{M}=128$. All inliers are quantized within a MaB and share the scale factor. Each MaB is then subdivided into multiple non-overlapping contiguous micro-blocks ($\mu$Bs) of size $B_{\mu}=8$ with sixteen $\mu$Bs forming a MaB. The outliers present in each $\mu$B shares same scale(s). As depicted in Figure 3(a), Step 1, the quantization process begins by first identifying inliers and outliers in each MaB by using the $3\sigma$ rule. A shared 8-bit power-of-two scale factor ($2^{I_{sf}}$), following Equation 1 is calculated for all inliers in a MaB and the inliers are quantized to 2-bit or 4-bit, resulting in MX-INT-(2/4)128 quantization. Interestingly, we observe that the inlier scale factor in each MaB is always a negative power of two for all FMs under consideration. We leverage this observation to reduce outlier magnitude, by multiplying all outlier values in a MaB with the inlier scale factor ($2^{I_{sf}}$) (this can also be perceived as division by $2^{-I_{sf}}$, for conformity with Equation 2). This pre-processing helps make outlier quantization easier, by pre-reducing its dynamic range before the actual outlier quantization.
Unlike inliers, outliers are quantized per $\mu$B, to reduce quantization error due to shared scaling over a larger group size (see section 7). After identifying outliers present in a $\mu$B, we compute a shared 8-bit MXScale that is calculated by concatenating the level-1 power-of-two scale factor ($2^{O^{l_{1}}_{sf}}$) and level-2 microExponent ($\mu X$). The level-1 scale factor is calculated by following Equation 1 to obtain 7 or 5-bit MSBs of MXScale depending on size of $\mu X$ being 1 or 3-bit of the LSBs–corresponds to exponent size of the FP format (depicted in Step 2 in Figure 3(a)). The outliers in a $\mu$B are scaled by ($2^{O^{l_{1}}_{sf}}$), following Equation 2 and then quantized to either e1m2/e3m4 FP-format (Rouhani et al., 2023) for $b$ = 4 or 8-bit, respectively. Post quantization of outliers, the level-two scale factor or the $\mu X$ is obtained by extracting the common exponent among all outliers in a $\mu$B. This process results in a MX-FP-$b$8,8 outlier quantization. The final outlier scale factor is $2^{O_{sf}}$ where, $O_{sf}$ is expressed as ${O_{sf}}={O^{l_{1}}_{sf}+\mu X-I_{sf}}$. The term $I_{sf}$ in the final outlier scale factor accounts for multiplication by $2^{I_{sf}}$ (or division by the inverse) during outlier pre-processing.